OpenStudy (anonymous):

Use the distributive property to simplify the expression below : -2(5d-1)+4d

OpenStudy (jiteshmeghwal9):

=-10d+2+4d =-6d+2.

OpenStudy (jackellyn):

=-2(5d) + (-2)(-1) +4d =-10d + 2 +4d =-10d +4d +2 =-6d +2